OpenStudy (anonymous):

I need to solve for x. Y=(T+6) divided by x Can someone help? I think that I multiply each side by (x/y) and end up with x=T+6 divided by Y. Is this right?

OpenStudy (anonymous):

\[y=\frac{t+6}{x}~~?\]

OpenStudy (anonymous):

If that's right, then yes, multiplying both sides by \(\dfrac{x}{y}\) will work: \[y\cdot\frac{x}{y}=\frac{t+6}{x}\cdot\frac{x}{y}~~\iff~~x=\frac{t+6}{y}\]

OpenStudy (anonymous):

Great! Thanks so much.